# Towards Fairness-Aware Adversarial Learning

### v1

@@ 1 Introduction
Deep learning models have undoubtedly achieved remarkable success across various domains, such as computer vision [30] and natural language processing [40]. However, they still remain susceptible to deliberate adversarial manipulations of input data [14, 45, 18, 19]. Robust learning techniques [28, 36, 21, 22] have emerged as a potential solution, aiming to enhance a model’s resilience against such vulnerabilities. These techniques have demonstrated a promising ability to enhance a model’s overall robustness, yet the intricate connection between robustness and fairness, as revealed by [42], demonstrates that the robust accuracy of the models can vary considerably across different categories or classes. Consider a scenario where an autonomous driving system attains commendable average robust accuracy in recognizing road objects. Despite this success, the system might demonstrate robustness against categories like inanimate objects (with high accuracy) while displaying vulnerability to crucial categories such as “human” (with low accuracy). This disparity or unfair robustness could potentially endanger drivers and pedestrians. Hence, it is vital to ensure consistent, equitable model performance against adversarial attacks by assessing worst-case robustness beyond average levels. This provides a more accurate evaluation than the average performance, recognizing the model’s limitations while ensuring reliability across diverse categories in real-world applications.
@@
Figure 1 provides an example, displaying the class-wise clean accuracy and robust accuracy (against AutoAttack [11]) using the Wide-ResNet34-10 [43] models on CIFAR-10 dataset, where both models are trained via standard training and adversarial training, respectively. It comes as no surprise that the model with standard training is vulnerable to adversarial attacks, yet it manages to attain comparable performance across various classes. In contrast, the adversarially trained model exhibits a noticeable bias, confidently classifying “automobile” but hesitating with “cat”. where robust accuracy diverges significantly across classes. Furthermore, it is noteworthy that even though the class “cat” attains the lowest clean and robust accuracy, the most significant disparity between clean and robust accuracy arises in the case of the class “deer”. This finding highlights an inconsistency between the clean and adversarial performances, particularly in terms of fairness where robust accuracy diverges significantly across classes. It is clear that the disparity between clean and robust accuracy is more pronounced in terms of fairness issues, where the robust accuracy on different classes illustrates severe diverges in the model.
@@
This phenomenon is called the “robust fairness” issue, which is first revealed in [42] referring to the gap between average robustness and worst-class robustness. Recently, some pioneering solutions have been proposed to address the robust fairness issue [34, 39, 25]. These efforts either tackle the robust fairness from the adversarial example generation or adjust the class weights empirically. However, essentially, most of these methods can be regarded as instances of reweighting, albeit diverse heuristic strategies are adopted. It is worth noting that such issues are absent in models trained without adversarial training, thus it is obvious that the underlying issue stems from the adversarial perturbations during the adversarial training. Inspired by but beyond these studies, we are prompted to consider why reweighting strategies are effective in mitigating the robust fairness issue. Intuitively, a model trained adversarially without reweighting fails to achieve high worst-class robust accuracy because it treats all classes equally, yet neglects the fact that the replaced adversarial examples may introduce bias to the final model. Conversely, reweighting can be perceived as a means of inducing a form of group distribution shift. This shift disrupts the uniform optimization of different classes, compelling the model to acquire resistance against these distribution shifts. This, in turn, leads to an enhancement in robust fairness. In line with this, we adopt an alternative approach to address this issue in the adversarial learning paradigm, where the following assumption is made:
@@
The robust fairness issue in the conventional AT is due to the unknown group (class) distribution shift induced by adversarial perturbations, which results in the overfitting problem.
@@
As opposed to the heuristic assignment for the reweighting item in the prior works, we believe that leveraging some optimization techniques for solving this overfitting problem can bring better results. Hence Distributional Robust Optimization (DRO) [3, 12] naturally emerges as a viable choice. Rather than assuming a fixed uniform data distribution, DRO acknowledges the inherent distributional uncertainty in real-world data, offering a more resilient and adaptable model structure. Therefore, this paper delves into the exploration and adaption of DRO, as a sensible solution for the robust fairness challenge. Specifically, after finding the adversarial example in adversarial training, instead of manually or empirically adjusting the weights for each class, we resort to learning the class-wise distributionally adversarial weights with the pre-defined constraints via DRO. By learning with these weights, the model will be guided to acquire the capacity to resist unknown group distribution shifts. The contributions of this paper are summarized as follows:
@@ 2.2 Distributional Robust Optimization
The origins of Distributional Robust Optimization (DRO) are found in the early studies on robust optimization [2, 7, 4], which eventually led to the development of DRO as a tool for handling distributional uncertainties. The application of DRO to machine learning problems has garnered significant attention like domain generalization [31], data distribution shift [32, 26], adversarial robustness [33, 8, 5] and traditional fairness in machine learning [15, 35, 23, 13]. While the intersection of DRO with robustness and fairness has begun to receive attention, there remain gaps in the literature, particularly in understanding how DRO can address robust fairness.
@@ 3.1 Preliminaries
When accounting for fairness, specifically the performance across different classes, Eq. 2 can be rewritten as:
It is noted that when a batch contains an equal number of data points for each class, Eq. 3 is technically identical to Eq. 2. According to the definition of distributional robust optimization [3, 12], we now consider minimizing the expected loss in the worst-case scenario over a set of uncertain distributions. This can be mathematically expressed as follows:
@@
To establish a connection between robust fairness and DRO, we can naturally delineate the classes as distinct groups within the training data. Subsequently, the uncertainty set can be defined with respect to these groups. Specifically, the setting of group DRO are borrowed [17, 29, 31], where the training distribution is assumed to be a mixture of groups (classes) indexed by . Thus the uncertainty set is defined as any mixture of these classes, i.e. , where is the probability simplex. Hence, the worst-case risk can be reformulated as the most detrimental combination across different groups, taking into account the expected loss for each individual class:
@@ 3.2 Problem definition
By disentangling the Eq. 5, it becomes evident that it can be interpreted as a reweighting objective of the ERM framework within the context of adversarial settings, incorporating the weighted factor . Empirical evidence has demonstrated the efficacy of several reweighting methods [42, 6, 39] in improving robust fairness. This aligns with the assumption stated in the introduction section, while we believe that leveraging optimization techniques, as opposed to heuristic assignment, will yield more effective advancements in promoting fairness.
As previously mentioned, the extensive range of distributions encompassed by the uncertainty set could present difficulties for DRO in sustaining its robustness. Nevertheless, the pivotal factor in addressing the group DRO lies in configuring the uncertainty set. To tackle this issue, we advocate an alternative approach to resolution: instead of relying solely on substantial regularization [31], we propose to use a straightforward yet effective ambiguity set with extra constraint. This is achieved by defining the ambiguity set as , where represents some distance metrics measuring the difference between two distributions, is the constraint parameter and is the uniform distribution. This choice of shrinks the width of the original and allows us to learn models that are robust to some group shifts, rather than identically uniform distribution among different classes. Therefore, our final objective in addressing the challenge posed by group distribution shifts within an adversarial setting can be expressed as follows:
@@ 3.3 Fairness-Aware Adversarial Learning
Based on the above objective, we propose a novel adversarial learning paradigm, named Fairness-Aware Adversarial Learning (FAAL)111Our code is available at https://github.com/TrustAI/FAAL, to improve the robust fairness via distributional robust optimization. Specifically, within the intermediary stage of the conventional adversarial training (between inner maximization and outer minimization), we introduce a class-wise distributionally adversarial weight for orientating the learning directions among different categories, which can be optimally solved by leveraging distributional robust optimization. By incorporating this weight in the outer minimization process to update the model’s parameters, the class (group) distribution shift can be protected, so the robust fairness issue will be alleviated.
To provide a clearer illustration of the comprehensive learning problem, we break down the entire learning task into three distinct stages:
@@
Phases 1 and 3 are the classic processes of conventional adversarial training, and Phase 2 is the core element of our proposed learning paradigm, as we assume that tackling the unknown class distributional shift can contribute to enhancing robust fairness. The pseudo-code of the whole learning framework can be seen in Algorithm 1. In the following content, we replace the notion of in Eq. 6 with for convenience and define it as Class-wise Distributionally Adversarial Weight.
@@
Given a class-wise objective loss on the adversarial examples, for all classes , the Class-wise Distributionally Adversarial Weight vector aims to maximize the overall loss:
@@
Given the loss in Definition 7 on the observed distribution, and suppose the regular loss on the test distribution with unknown group distribution shift, then the following holds for all :
@@
Theorem 1 tells that the is guaranteed to be the upper bound of with high probability. In line with this, it enjoys a strong generalization where the performance on the test distribution with some unknown group distribution shift is at least as good as the estimated performance with high probability. So the solution of the distributionally adversarial weight solving by the convex optimization is optimal and will provide protection on the unknown class distribution shift. As cross-entropy loss cannot well-represents how good the discrepancy between classes [9], we instead use the CW margin loss [10] as for calculating the class-wise distributionally adversarial weight:
@@
| + Fine-tune with FRL-RWRM | 80 | 83.25 (74.80) | 50.37 (38.10) | 49.77 (36.60) | 46.97 (33.10) |
| + Fine-tune with FRL-RWRM | 80 | 85.12 (71.60) | 52.56 (37.10) | 51.92 (35.50) | 49.60 (31.70) |
@@
| + Fine-tune with FRL-RWRM | 80 | 82.90 (72.70) | 53.16 (40.60) | 51.39 (36.30) | 49.97 (35.40) |
| + Fine-tune with FRL-RWRM | 80 | 85.19 (70.90) | 53.76 (39.20) | 52.92 (36.80) | 51.30 (34.60) |
@@
| + Fine-tune with FRL-RWRM | 80 | 83.72 (71.80) | 52.16 (37.50) | 50.73 (35.00) | 49.19 (31.70) |
| + Fine-tune with FRL-RWRM | 80 | 82.09 (71.80) | 50.86 (36.00) | 49.78 (33.00) | 47.78 (30.30) |
@@
| + Fine-tune with FRL-RWRM | 80 | 82.31 (65.90) | 49.90 (31.70) | 49.68 (34.00) | 46.50 (27.70) |
| + Fine-tune with FRL-RWRM | 80 | 84.24 (65.70) | 48.63 (30.90) | 49.77 (31.50) | 46.53 (28.60) |
@@ 4 Experimental Results
Given our method’s focus on the robust fairness challenge, it is reasonable to assume that the model already possesses a certain degree of average robustness. Otherwise, considering the robust fairness issue might not yield meaningful results. Hence, the question arises: Is it imperative to initiate the training of a model from the beginning for achieving fairness with a certain robustness level? In the next section, we first test our approach through adversarial fine-tuning, and then explore if training from scratch with our method offers any additional benefits.
@@ 4.1 Fine-tuning for Enhancing Robust Fairness
Baselines and experiment settings: We first conducted experiments on CIFAR-10 dataset [24], which is popularly used for adversarial training evaluation. We use the average & worst-class accuracy under different adversarial attacks (Clean / PGD [28] / CW [10] / AutoAttack [11]) as the evaluation metrics. The perturbation budget is set to on CIFAR-10 dataset. FRL [42] is the only existing state-of-the-art technique from the recent literature which performs fine-tuning to a pre-trained model for improving robust fairness. FRL proposed two strategies based on TRADES [44] for enhancing robust fairness, including reweight (RW) and remargin (RM). Hence we apply the best versions of FRL from their paper: FRL-RWRM with and FRL-RWRM with , where and are the fairness constraint parameters for reweight and remargin of FRL, we name them FRL-RWRM and FRL-RWRM for short. The results of FRL are reproduced using their public code, where the target models are fine-tuned for 80 epochs and the best results are presented. In terms of the proposed method, although we utilize the PGD-AT adversary method by default (named FAAL for short), it is completely compatible with other AT approaches. We found that 2 epochs of fine-tuning are enough to improve the robust fairness greatly, without sacrificing too much average clean/robust accuracy. We set the value of in our method as , and the learning rate is configured from in the first epoch and drops to in the second epoch.
@@
Table 1 demonstrates our main results of finetuning Wide-Resnet34-10 (WRN-34-10) models [43] on CIFAR-10 dataset, where different state-of-the-art adversarial defensed methods are adopted, including PGD-AT [28], TRADES [44], MART [38] and AWP [41]. We can see that FAAL outperforms the two FRL methods with respect to both average and worst-class robustness. Notably, in the majority of cases, FAAL achieves this without significant compromises in clean accuracy, unlike the FRL methods which tend to trade off the clean accuracy for improving robustness. FAAL promotes the worst-class AutoAttack (AA) accuracy by approximately than FRL for fine-tuning PGD-AT, MART, and AWP. Except for fine-tuning TRADES, both methods yield comparable performance, this is due to that FRL is a TRADES-based method and it takes advantage of knowing the source method. Most importantly, FRL requires many epochs (80 epochs) to obtain the best results, while our method, is able to achieve better results within only 2 epochs. As adversarially training a large model with high robustness is already time-consuming, to circumvent the need for retraining the model from the beginning, FAAL offers a solution for saving time and computational resources. It demonstrates the capability to quickly fine-tune a robust model that initially lacks fairness, resulting in a model that is both robust and fair. Due to the space limit, similar improvements on the Preact-Resnet model can be found in the Appendix.
@@
Strong adversarial attacks can help? The remargin of FRL [42] claims that increasing the perturbation margin can help for obtaining better robust fairness, while this may hurt the average clean accuracy/robustness, as indicated in Tab. 1. Certainly, there is a trade-off existing between the average robustness and worst-class robustness, but is it necessary to increase the perturbation margin for improving the class-wise robustness? We question whether this is a mandatory requirement for improvements, and we assume the benefits come from the stronger strength of adversarial perturbation. Hence, instead of enlarging the perturbation margin, we capitalize on the flexibility of our learning framework and integrate our method with AWP [41], a well-regarded model weight perturbation technique, to strength the attacks. As shown in Tab. 1, when combining with AWP, FAAL further enhances the worst-class robust accuracy on WRN34-10 models, especially for the original one is adversarial trained with AWP. FAAL is almost unharmed on the improvement to the original unfair models most of the time. Therefore, it is not compulsory to enlarge the perturbation margin to gain better results, where applying a stronger adversary indeed benefits robust fairness without enlarging the perturbation margin. Figure 2 visualizes the results of class-wise AA accuracy for the comparison of the proposed method FAAL and FAAL, and two FRL baselines, respectively. It can be seen that FAAL clearly boost the worst-class robust accuracy, presenting outstanding capacity to improve robust fairness with high effectiveness and efficiency, respectively, where it outperforms FRL not only for the average/worst-class robustness but also for the very rare fine-tuning steps.
@@
| CFA | 80.82 (64.60) | 50.10 (24.40) |
| CFA | 80.36 (66.20) | 50.10 (26.50) |
| WAT | 80.37 (66.00) | 46.16 (30.70) |
| FAAL | 82.20 (62.90) | 49.10 (33.70) |
| FAAL | 81.62 (68.90) | 48.48 (33.60) |

### v2

@@ 1 Introduction
Deep learning models have undoubtedly achieved remarkable success across various domains, such as computer vision [31, 46] and natural language processing [41]. However, they still remain susceptible to deliberate adversarial manipulations of input data [15, 48, 49, 19, 20, 44]. Adversarial training techniques [29, 37, 22, 23, 11] have emerged as a potential solution, aiming to enhance a model’s resilience against such vulnerabilities. These techniques have demonstrated a promising ability to enhance a model’s overall robustness, yet the intricate connection between robustness and fairness, as revealed by [43], demonstrates that the robust accuracy of the models can vary considerably across different categories or classes. Consider a scenario where an autonomous driving system attains commendable average robust accuracy in recognizing road objects; despite this success, the system might demonstrate robustness against categories like inanimate objects (with high accuracy) while displaying vulnerability to crucial categories such as “human" (with low accuracy). This disparity or unfair robustness could potentially endanger drivers and pedestrians. Hence, it is vital to ensure consistent, equitable model performance against adversarial attacks by assessing worst-case robustness beyond average levels. This provides a more accurate evaluation than the average performance, recognizing the model’s limitations while ensuring reliability across diverse categories in real-world applications.
@@
Figure 1 provides an example, displaying the class-wise clean accuracy and robust accuracy against AutoAttack [12] using the Wide-ResNet34-10 [45] models on CIFAR-10 dataset, where both models are trained via standard training and adversarial training, respectively. It comes as no surprise that the model with standard training is vulnerable to adversarial attacks, yet it manages to attain comparable performance across various classes. In contrast, the adversarially trained model exhibits a noticeable bias, confidently classifying “automobile" but hesitating with “cat". That is, the robust accuracy diverges significantly across classes. Furthermore, it is noteworthy that even though the class “cat" attains the lowest clean and robust accuracy, the most significant disparity between clean and robust accuracy arises in the case of the class “deer". This finding highlights an inconsistency between the clean and adversarial performances, where the robust accuracy on different classes illustrates more severe diverges in the model.
@@
This phenomenon is called the “robust fairness" issue, which is first revealed in [43] referring to the gap between average robustness and worst-class robustness. Recently, some pioneering solutions have been proposed to address the robust fairness issue [35, 40, 26]. These efforts either tackle the robust fairness from the adversarial example generation or adjust the class weights empirically. However, essentially, most of these methods can be regarded as instances of reweighting, albeit diverse heuristic strategies are adopted. It is worth noting that such issues are absent in models trained without adversarial training, thus it is obvious that the underlying issue stems from the adversarial perturbations during the adversarial training. Inspired by but beyond these studies, we are prompted to consider why reweighting strategies are effective in mitigating the robust fairness issue. Intuitively, a model trained adversarially without reweighting fails to achieve high worst-class robust accuracy because it treats all classes equally, yet neglects the fact that the replaced adversarial examples may introduce bias to the final model. Conversely, reweighting can be perceived as a means of inducing a form of group distribution shift. This shift disrupts the uniform optimization of different classes, compelling the model to acquire resistance against these distribution shifts. This, in turn, leads to an enhancement in robust fairness. In line with this, we adopt an alternative approach to address this issue in the adversarial learning paradigm, where the following assumption is made:
@@
The robust fairness issue in the conventional AT is due to the unknown group (class) distribution shift induced by the generated adversarial perturbations, which results in the overfitting problem.
@@
As opposed to the heuristic assignment for the reweighting item in the prior works, instead, we are expecting to leverage some optimization techniques for reweighting, such that may bring better results to resolve this overfitting problem. Hence Distributional Robust Optimization (DRO) [3, 13] naturally emerges as a viable choice. Rather than assuming a fixed uniform data distribution, DRO acknowledges the inherent distributional uncertainty in real-world data, offering a more resilient and adaptable model structure. Therefore, this paper delves into the exploration and adaption of DRO, as a sensible solution for the robust fairness challenge. Specifically, after finding the adversarial example in adversarial training, instead of manually or empirically adjusting the weights for each class, we resort to learning the class-wise distributionally adversarial weights with the pre-defined constraints via DRO. By learning with these weights, the model will be guided to acquire the capacity to resist unknown group distribution shifts. The contributions of this paper are summarized as follows:
@@ 2.2 Distributional Robust Optimization
The origins of DRO are found in the early studies on robust optimization [2, 7, 4], which eventually led to the development of DRO as a tool for handling distributional uncertainties. The application of DRO to machine learning problems has garnered significant attention like domain generalization [32], data distribution shift [33, 27], adversarial robustness [34, 8, 5] and traditional fairness in machine learning [16, 36, 24, 14]. While the intersection of DRO with robustness and fairness has begun to receive attention, there remain gaps in the literature, particularly in understanding how DRO can address this fairness in an adversarial setting.
@@ 3.1 Preliminaries
When accounting for class fairness, specifically the performance across different classes, Eq. 2 can be rewritten as:
It is noted that when a batch contains an equal number of data points for each class, Eq. 3 is technically identical to Eq. 2. According to the definition of distributional robust optimization [3, 13], we now consider minimizing the expected loss in the worst-case scenario over a set of uncertain distributions. This can be mathematically expressed as:
@@
To establish the connection between robust fairness and DRO, we can naturally delineate the classes as distinct groups within the training data. Subsequently, the uncertainty set can be defined with respect to these groups. Specifically, the setting of group DRO are borrowed [18, 30, 32], where the training distribution is assumed to be a mixture of groups (classes) indexed by . Thus the uncertainty set is defined as any mixture of these classes, i.e. , where is the probability simplex. Hence, the worst-case risk can be reformulated as the most detrimental combination across different groups, taking into account the expected loss for each class:
@@ 3.2 Problem Definition
By disentangling the Eq. 5, it becomes evident that it can be interpreted as a reweighting objective of the ERM framework within the context of adversarial settings, incorporating with the weighted factor . Empirical evidence has demonstrated the efficacy of several reweighting methods [43, 6, 40] in improving robust fairness and all of them fall into the same paradigm of Eq. 5. This aligns with the assumption stated in the introduction section, while we seek to leverage some optimization techniques for promoting fairness directly, rather than a heuristic assignment.
As previously mentioned, the extensive range of distributions encompassed by the uncertainty set could present difficulties for DRO in sustaining its robustness. Nevertheless, the pivotal factor in addressing the group DRO lies in configuring the uncertainty set. To tackle this issue, we advocate an alternative solution: instead of relying solely on substantial regularization [32], we propose to use a straightforward yet effective ambiguity set with extra constraint. This is achieved by defining the ambiguity set as , where represents some distance metrics measuring the difference between two distributions, is the constraint parameter and is the uniform distribution. This choice of shrinks the width of and allows us to learn models that are robust to some group shifts, rather than identically uniform distribution among different classes. Hence, our final objective in addressing the challenge posed by potential group distribution shifts within an adversarial setting can be denoted as:
@@ 3.3 Fairness-Aware Adversarial Learning
Based on the above objective, we propose a novel adversarial learning paradigm, named Fairness-Aware Adversarial Learning (FAAL)111Our code is available at https://github.com/TrustAI/FAAL, to improve the robust fairness via distributional robust optimization. Specifically, within the intermediary stage of the conventional adversarial training, i.e. between inner maximization and outer minimization, we introduce a class-wise distributionally adversarial weight for orientating the learning directions among different categories, which can be optimally solved by leveraging distributional robust optimization. By incorporating this weight into the outer minimization process to update the model’s parameters, the class (group) distribution shift can be protected to alleviate the robust fairness issue.
To provide a clearer illustration of the whole learning problem, we break down it into three distinct stages:
@@
Phases 1 and 3 are the classic processes of conventional adversarial training, and Phase 2 is the core element of our proposed learning paradigm, as we assume that tackling the unknown class distributional shift can contribute to enhancing robust fairness. The whole procedure is summarized in Algorithm 1. In the following content, we replace the notion of in Eq. 6 with for convenience and define it as Class-wise Distributionally Adversarial Weight.
@@
Given a class-wise objective loss on the adversarial examples, for all classes , the optimal Class-wise Distributionally Adversarial Weight vector aims to maximize the overall loss:
@@
Given the loss in Eq. 7 on the observed distribution, and suppose the regular loss on the test distribution with unknown group distribution shift, then the following holds for all :
@@
Theorem 1 tells that the is guaranteed to be the upper bound of with high probability given the large number of observed sample . In line with this, it enjoys a strong generalization where the performance on the test distribution with some unknown group distribution shift is at least as good as the estimated performance with high probability. So the solution of the class-wise distributionally adversarial weight solving by the convex optimization is optimal and will provide protection on the unknown class distribution shift. As cross-entropy loss cannot well-represents how good the discrepancy between classes [9], we instead use the CW margin loss [10] as for calculating the class-wise distributionally adversarial weight:
@@
| + Fine-tune with FRL-RWRM0.05 | 80 | 83.25 (74.80) | 50.37 (38.10) | 49.77 (36.60) | 46.97 (33.10) |
| + Fine-tune with FRL-RWRM0.07 | 80 | 85.12 (71.60) | 52.56 (37.10) | 51.92 (35.50) | 49.60 (31.70) |
@@
| + Fine-tune with FRL-RWRM0.05 | 80 | 82.90 (72.70) | 53.16 (40.60) | 51.39 (36.30) | 49.97 (35.40) |
| + Fine-tune with FRL-RWRM0.07 | 80 | 85.19 (70.90) | 53.76 (39.20) | 52.92 (36.80) | 51.30 (34.60) |
@@
| + Fine-tune with FRL-RWRM0.05 | 80 | 83.72 (71.80) | 52.16 (37.50) | 50.73 (35.00) | 49.19 (31.70) |
| + Fine-tune with FRL-RWRM0.07 | 80 | 82.09 (71.80) | 50.86 (36.00) | 49.78 (33.00) | 47.78 (30.30) |
@@
| + Fine-tune with FRL-RWRM0.05 | 80 | 82.31 (65.90) | 49.90 (31.70) | 49.68 (34.00) | 46.50 (27.70) |
| + Fine-tune with FRL-RWRM0.07 | 80 | 84.24 (65.70) | 48.63 (30.90) | 49.77 (31.50) | 46.53 (28.60) |
@@ 4 Experimental Results
Given our method’s focus on the robust fairness challenge, it is reasonable to assume that the model already possesses a certain degree of average robustness. Otherwise, considering the robust fairness issue might not yield meaningful results. Hence, the question arises: Is it imperative to initiate the training of a model from the beginning for achieving fairness with a certain robustness level? In the next section, we first test our approach through adversarial fine-tuning, and then explore if training from scratch with our method can offer additional benefits.
@@ 4.1 Fine-tuning for Enhancing Robust Fairness
Baselines and experiment settings: We first conducted experiments on CIFAR-10 dataset [25], which is popularly used for adversarial training evaluation. We use the average & worst-class accuracy under different adversarial attacks (Clean / PGD [29] / CW [10] / AutoAttack [12]) as the evaluation metrics. The perturbation budget is set to on CIFAR-10 dataset. FRL [43] is the only existing state-of-the-art technique from the recent literature which performs fine-tuning to a pre-trained model for improving robust fairness. FRL proposed two strategies based on TRADES [47] for enhancing robust fairness, including reweight (RW) and remargin (RM). Hence we apply the best versions of FRL from their paper: FRL-RWRM with and FRL-RWRM with , where and are the fairness constraint parameters for reweight and remargin of FRL, we name them FRL-RWRM0.05 and FRL-RWRM0.07 for short. The results of FRL are reproduced using their public code, where the target models are fine-tuned for 80 epochs and the best results are presented.
In terms of the proposed method, although we utilize the PGD-AT adversary method by default (named FAALAT for short), it is completely compatible with other AT approaches like TRADES or MART. To achieve this, one just needs to keep the original implementation for both inner maximization and outer minimization unchanged and add the intermediate maximization independently. We found that 2 epochs of fine-tuning are enough to improve the robust fairness greatly, without sacrificing too much average clean/robust accuracy. We set the value of in our method as , and the learning rate is configured from in the first epoch and drops to in the second epoch.
@@
Table 1 demonstrates our main results of finetuning Wide-Resnet34-10 (WRN-34-10) models [45] on CIFAR-10 dataset, where different state-of-the-art adversarial defensed methods are adopted, including PGD-AT [29], TRADES [47], MART [39] and AWP [42]. We can see that FAALAT outperforms the two FRL methods with respect to both average and worst-class robustness. Notably, in the majority of cases, FAALAT achieves this without significant compromises in clean accuracy, unlike FRL methods which tend to trade off the clean accuracy for improving robustness. FAALAT promotes the worst-class AutoAttack (AA) accuracy by approximately than FRL for fine-tuning PGD-AT, MART, and AWP. Except for fine-tuning TRADES, both methods yield comparable performance, this is partially due to that FRL is a TRADES-based method and it takes advantage of knowing the source method. Most importantly, FRL requires many epochs (80 epochs) to obtain the best results, while our method, is able to achieve better results within only 2 epochs. As adversarially training a large model with high robustness is already time-consuming, to circumvent the need for retraining the model from the beginning, FAAL offers a solution for saving time and computational resources. It demonstrates the capability to quickly fine-tune a robust model that initially lacks fairness, resulting in a model that is both robust and fair. Due to the space limit, similar improvements on the Preact-Resnet model can be found in the Appendix.
@@
Strong adversarial attacks can help? The remargin of FRL [43] claims that increasing the perturbation margin can help for obtaining better robust fairness, while this may hurt the average clean accuracy/robustness, as indicated in Tab. 1. Certainly, there is a trade-off existing between the average robustness and worst-class robustness, but is it necessary to increase the perturbation margin for improving the class-wise robustness? We question whether this is a mandatory requirement for improvements, and we assume the benefits come from the stronger strength of adversarial perturbation. Hence, instead of enlarging the perturbation margin, we capitalize on the flexibility of our learning framework and integrate our method with AWP [42], a well-regarded model weight perturbation technique, to strengthen the attacks. As shown in Tab. 1, when combining with AWP, FAALAT-AWP further enhances the worst-class robust accuracy on WRN34-10 models, especially for the original one adversarially trained with AWP. FAALAT-AWP is almost unharmed on the improvement to the original unfair models most of the time. Therefore, it is not compulsory to enlarge the perturbation margin to gain better results, where applying a stronger adversary indeed benefits robust fairness without enlarging the perturbation margin. Figure 2 visualizes the results of class-wise AA accuracy for the comparison of the proposed method FAALAT and FAALAT-AWP, and two FRL baselines, respectively. It can be seen that FAAL boost the worst-class robust accuracy, presenting outstanding capacity to improve robust fairness with high effectiveness and efficiency, respectively, where it outperforms FRL not only for the average/worst-class robustness but also for the very rare fine-tuning steps.
@@
| CFAAT | 80.82 (64.60) | 50.10 (24.40) |
| CFATRADES | 80.36 (66.20) | 50.10 (26.50) |
| WATTRADES | 80.37 (66.00) | 46.16 (30.70) |
| FAALAT | 82.20 (62.90) | 49.10 (33.70) |
| FAALTRADES | 81.62 (68.90) | 48.48 (33.60) |